write the equation of a parabola that passes through (1,-1) and has a vertex at (-3,3)
Answer by lwsshak3(11628)

write the equation of a parabola that passes through (1,-1) and has a vertex at (-3,3)
Standard form of equation for a parabola with vertex (-3,3):
y=A(x-h)^2+k, (h,k)=vertex
Solving for A(Using given coordinates (1,-1)passing through and (-3,3 of vertex.
-1=A(1+3)^2+3
0=A(4)^2+4
16A=-4
A=-4/16=-1/4
Equation:y=-1/4(x+3)^2+3